# Laypeople’s Egocentric Perceptions of Copyright for AI-Generated Art

### v1

@@ 5.2.3. Attribution of Rights
Finally, we also identified some differences between treatments regarding the right to display AI-generated images: invested evaluators were less likely to support non-exclusive rights for creators’ images than uninvested evaluators and the creators themselves. A potential explanation is that creators would like audiences to see their creations because it could lead to further financial benefits in the long run. Instead of keeping their creations “behind curtains,” creators could have imagined that allowing others to display their images could bring financial benefits down the road. Furthermore, invested evaluators might have wanted to “hide” their competitors’ creations because it competes with their own submissions, while uninvested evaluators had no incentive to do so. Additional longitudinal studies are necessary to explore this hypothesis further.
@@
Participants evaluated their own creations egocentrically when this assessment was supposed to determine who received monetary awards. On average, creators rated their images over 1.8 point higher (on a 11-point scale) than evaluators evaluated the same images. Furthermore, invested evaluators evaluated their own images around 1.5 point higher than they evaluated creators’ images. Our results suggest that financial incentives can bias people’s responses to AI-generated art. However, when asked who should hypothetically have some of the rights associated with copyright protection, participants did not prioritize their own gains. Instead, egocentric biases emerged only in the context of actual rewards, and not hypothetical rights, suggesting that future debates concerning AI-generated works in relation to copyright law may be tainted by conflicts of interest.
Participants’ tendency to rate their creations egocentrically when financial incentives are present but not when asked concerning hypothetical rights is also relevant in relation to data contributors. We found that people acknowledged data contributors as authors and potential rights-holders; yet, it is unclear whether we would observe the same result had the corresponding questions involved actual monetary awards. It is possible that if participants were asked to financially compensate data contributors, e.g., by sharing their award, people would not do it and rather keep the incentive to themselves. Such a finding would suggest that laypeople may support regulatory frameworks that compensate data contributors only if these frameworks do not affect them financially. Our study design could be easily modified to address these research questions. For instance, future studies could include questions concerning participants’ willingness to share their potential reward with data contributors.
It is worth noting that while creators and invested evaluators assigned higher ratings to their own images, we found no evidence of the sabotage hypothesized in H3. That is, invested evaluators did not give creators significantly lower scores than uninvested evaluators. It is unclear whether this lack of sabotage is the result of a lack of understanding of the incentive structure or if it reflects participants true altruistic behavior. Future research is needed to test participants’ understanding of the incentive structure. Another potential explanation is that participants chose to inflate their own scores instead of lowering other participants’ scores because doing so could be seen less morally contentious—a hypothesis that future work could explore further.
@@ 5.3. Concluding Remarks
In summary, we found that users of GenAI judge their own creations egocentrically concerning some factors, but not others. In fact, our results even suggest effects in the opposite direction for some of the variables we examined (e.g., perceived authorship of the user). Importantly, we identified the importance of financial interests in determining whether people favor their own AI-generated images compared to other people’s creations.
@@ Acknowledgements.
- Miller and Ross (1975) Dale T Miller and Michael Ross. 1975. Self-serving biases in the attribution of causality: Fact or fiction? Psychological bulletin 82, 2 (1975), 213.
@@
- Shepperd et al. (2008) James Shepperd, Wendi Malone, and Kate Sweeny. 2008. Exploring causes of the self-serving bias. Social and Personality Psychology Compass 2, 2 (2008), 895–908.

### v2

@@ 5.2.3. Attribution of Rights
Finally, we also identified some differences between treatments regarding the right to display AI-generated images: creators were more likely to support non-exclusive rights for their own images than invested evaluators evaluating the same images. A potential explanation is that creators would like audiences to see their creations because it could lead to further financial benefits in the long run. Instead of keeping their creations “behind curtains,” creators could have imagined that allowing others to display their images could bring financial benefits down the road. Additional longitudinal studies are necessary to explore this hypothesis further.
@@
Participants evaluated their own creations egocentrically when this assessment was supposed to determine who received monetary awards. On average, creators rated their images approximately 1.9 points higher (on an 11-point scale) than evaluators evaluating the same images. Furthermore, invested evaluators evaluated their own images around 1.5 points higher than they evaluated creators’ images. Our results suggest that financial incentives can bias people’s responses to AI-generated art. However, when asked who should hypothetically have some of the rights associated with copyright protection, participants did not prioritize their own gains. Instead, egocentric biases emerged only in the context of actual rewards, and not hypothetical rights, suggesting that future debates concerning AI-generated works in relation to copyright law may be tainted by conflicts of interest.
Participants’ tendency to rate their creations egocentrically when financial incentives are present but not when asked about hypothetical rights is also relevant in relation to data contributors. We found that people acknowledged data contributors as authors and potential rights-holders; yet, it is unclear whether we would observe the same result had the corresponding questions involved actual monetary awards. It is possible that if participants were asked to financially compensate data contributors, e.g., by sharing their award, people would not do it and would rather keep the incentive to themselves. Such a finding would suggest that laypeople may support regulatory frameworks that compensate data contributors only if these frameworks do not affect them financially. Our study design could be easily modified to address these research questions. For instance, future studies could include questions concerning participants’ willingness to share their potential reward with data contributors.
It is worth noting that while creators and invested evaluators assigned higher ratings to their own images, we found no evidence of the sabotage hypothesized in H3. That is, invested evaluators did not give creators significantly lower scores than uninvested evaluators. It is unclear whether this lack of sabotage is the result of a lack of understanding of the incentive structure or if it reflects participants’ true altruistic behavior. Future research is needed to test participants’ understanding of the incentive structure. Another potential explanation is that participants chose to inflate their own scores instead of lowering other participants’ scores because doing so could be seen as less morally contentious—a hypothesis that future work could explore further.
@@
The fact that the potential of monetary awards changed people’s perceptions of GenAI outputs also has implications for the design of AI models. There are already examples of how AI-generated art could give users financial benefits (Roose, 2022) even though its legal status is still undefined. Future work could explore design patterns for GenAI products that are more effectively transparent about potential legal and financial consequences so that users are aware of both positive and negative outcomes (as discussed above in relation to perceived authorship).
@@ 5.3. Concluding Remarks
In summary, we found that users of GenAI judge their own creations egocentrically concerning some factors but not others. In fact, our results even suggest effects in the opposite direction for some of the variables we examined (e.g., perceived authorship of the user). Importantly, we identified the importance of financial interests in determining whether people favor their own AI-generated images compared to other people’s creations.
@@ Acknowledgements.
- Amazon (2024) Amazon. 2024. Amazon announces AI Shopping Guides, generative AI-powered buying guides that seamlessly bring together product research and relevant products. https://tinyurl.com/3epksaee.
@@
- Beyond the Veil Art ([n. d.]) Beyond the Veil Art. [n. d.]. What is an art exhibition? https://www.beyondtheveilart.com/faq/what-is-an-art-exhibition/. (Accessed on 08/27/2024).
@@
- Calacci and Pentland (2022) Dan Calacci and Alex Pentland. 2022. Bargaining with the black-box: Designing and deploying worker-centric tools to audit algorithmic management. Proceedings of the ACM on Human-Computer Interaction 6, CSCW2 (2022), 1–24.
@@
- Feier et al. (2022) Till Feier, Jan Gogoll, and Matthias Uhl. 2022. Hiding behind machines: Artificial agents may help to evade punishment. Science and Engineering Ethics 28, 2 (2022), 19.
- Fiesler (2020) Casey Fiesler. 2020. Lawful users: Copyright circumvention and legal constraints on technology use. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems. 1–11.
@@
- Fraser (2024) Colin Fraser. 2024. Generative AI is a hammer and no one knows what is and isn’t a nail. Medium. https://tinyurl.com/teybczjx.
- Garrett et al. (2020) Natalie Garrett, Nathan Beard, and Casey Fiesler. 2020. More than” If Time Allows” the role of ethics in AI education. In Proceedings of the AAAI/ACM Conference on AI, Ethics, and Society. 272–278.
@@
- Goetze (2024) Trystan S Goetze. 2024. AI Art is Theft: Labour, Extraction, and Exploitation: Or, On the Dangers of Stochastic Pollocks. In The 2024 ACM Conference on Fairness, Accountability, and Transparency. 186–196.
@@
- Gulotta et al. (2013) Rebecca Gulotta, William Odom, Jodi Forlizzi, and Haakon Faste. 2013. Digital artifacts as legacy: exploring the lifespan and value of digital data. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems. 1813–1822.
@@
- Köbis et al. (2021) Nils Köbis, Jean-François Bonnefon, and Iyad Rahwan. 2021. Bad machines corrupt good morals. Nature human behaviour 5, 6 (2021), 679–685.
@@
- Krafft et al. (2021) PM Krafft, Meg Young, Michael Katell, Jennifer E Lee, Shankar Narayan, Micah Epstein, Dharma Dailey, Bernease Herman, Aaron Tam, Vivian Guetler, et al. 2021. An action-oriented AI policy toolkit for technology audits by community advocates and activists. In Proceedings of the 2021 ACM conference on fairness, accountability, and transparency. 772–781.
@@
- Lazar et al. (2019) Jonathan Lazar, Raja Kushalnagar, Casey Fiesler, and G Anthony Giannoumis. 2019. Introduction to Legal Issues in Human-Computer Interaction. In Extended Abstracts of the 2019 CHI Conference on Human Factors in Computing Systems. 1–5.
@@
- Midjourney (2024) Midjourney. 2024. Terms of Service. https://tinyurl.com/4rrk28xt.
@@
- Mink et al. (2022) Jaron Mink, Licheng Luo, Natã M Barbosa, Olivia Figueira, Yang Wang, and Gang Wang. 2022. DeepPhish: Understanding user trust towards artificially generated profiles in online social networks. In 31st USENIX Security Symposium (USENIX Security 22). 1669–1686.
- Mink et al. (2024) Jaron Mink, Miranda Wei, Collins W Munyendo, Kurt Hugenberg, Tadayoshi Kohno, Elissa M Redmiles, and Gang Wang. 2024. It’s Trying Too Hard To Look Real: Deepfake Moderation Mistakes and Identity-Based Bias. In Proceedings of the CHI Conference on Human Factors in Computing Systems. 1–20.
@@
- OpenAI (2024) OpenAI. 2024. Terms of use. https://tinyurl.com/bdwbtc8m.
@@
- Reid (2024) Elizabeth Reid. 2024. Supercharging Search with generative AI. Google. https://tinyurl.com/ywdnkn8n.
@@
- Stability.AI (2022) Stability.AI. 2022. Stability AI API Terms of Service. https://tinyurl.com/f4u6dmww.
@@
- WhatsAp (2024) WhatsAp. 2024. How to create and share AI stickers. WhatsApp FAQ. https://tinyurl.com/4f3wtsth.
- Widder et al. (2024) David Gray Widder, Laura Dabbish, James D Herbsleb, and Nikolas Martelaro. 2024. Power and Play: Investigating” License to Critique” in Teams’ AI Ethics Discussions. Proceedings of the ACM on Human-Computer Interaction 8, CSCW2 (2024), 1–23.
- Widder et al. (2023) David Gray Widder, Derrick Zhen, Laura Dabbish, and James Herbsleb. 2023. It’s about power: What ethical concerns do software engineers have, and what do they (feel they can) do about them?. In Proceedings of the 2023 ACM Conference on Fairness, Accountability, and Transparency. 467–479.
- Xu and Warschauer (2020) Ying Xu and Mark Warschauer. 2020. What are you talking to?: Understanding children’s perceptions of conversational agents. In Proceedings of the 2020 CHI conference on human factors in computing systems. 1–13.